Evaluate The integrals.

5. Integral of t^2 + 3t + 1/t^2 dt
6. Integral of (2x^3-1)(1/ x^2 + 5)
1
Expert's answer
2019-11-06T10:33:05-0500

"5.\smallint" (t2+3t+1/t2)=t3/3+3t2/2-1/t+C;

6. "\smallint" (2x3-1)(1 /x2+5)dx="\int" (2x+10x3-1/x2-5)dx=2x2/2+10x4/4+1/x-5x +C